# On Utility Maximization under Multivariate Fake Stationary Affine Volterra Models

### 11 Mar 2026 v1

@@ 5 Proofs of the main results
${Z}_{t}:=\mathcal{E}(\int_{0}^{t}g(s)\alpha_{s}^{\top}\,dB_{s})=\exp\Big(\int_{0}^{t}g(s)\alpha_{s}^{\top}\,dB_{s}-\frac{1}{2}\int_{0}^{t}g^{2}(s)\left|\alpha_{s}\right|^{2}ds\Big),\quad t\in[0,T].$ (5.82)
@@ 2. $(b)$
We provide a proof of this classical result for the reader’s convenience.
Proof: The first claim in (5.85) follows by the associativity of the convolution and applying the fundamental theorem of calculus. For the second claim, we could use calculus of convolutions and resolvents. However, we rather used Laplace transform and deduce that:
5.1 Proofs of Proposition 3.4 and Theorems 3.5 and 3.7
Proof of Proposition 3.4 : We write for  $0\leq t\leq T$:
$\Gamma_{t}^{\frac{1}{\delta}}=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}\frac{\gamma}{\delta}\big(r(s)+\frac{\left|\lambda_{s}\right|^{2}}{2(1-\gamma)}\big)ds\Big)\mid\mathcal{F}_{t}\Big]=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}\frac{\gamma}{\delta}\big(r(s)+\frac{1}{2(1-\gamma)}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}\big)ds\Big)\mid\mathcal{F}_{t}\Big]$ (5.87)
which ensures that $\Gamma_{t}>0$ $\mathbb{P}-a.s.$, since $V_{t}$ is non-negative ($V\in\mathbb{R}^{d}_{+}$), $r(t)>0$ is deterministic, and $1-\gamma\leq\delta\leq 1$. We then have in view of (5.87) that there exists some positive constant $m>0$ such that $\Gamma_{t}\geq m>0$ for every $t\in[0,T]$. An application of the exponential-affine transform formula (Gnabeyeu, 2026, Theorem A.4.) with $\mathcal{M}\ni m(\,\mathrm{d}s):=\frac{\gamma}{2\delta(1-\gamma)}\theta\odot\theta\,{\rm Leb}_{d}(\,\mathrm{d}s)$ (where $\odot$ denote the Hadamard (pointwise or component-wise) product) yields:
[…]
where we used for the last identity the fact that
where the drift factor takes the form:
$\hskip 142.26378pt\forall\,t\in[0,T]\quad J_{t}^{\alpha}=\frac{\Gamma_{0}x_{0}^{\gamma}}{\gamma}e^{\int_{0}^{t}D_{s}(\alpha_{s})\,ds}F_{t}^{\alpha}\quad\text{where}$ (5.88)
@@
$X_{t}^{\alpha^{*}}=x_{0}\exp\!\left(\int_{0}^{t}\Big(r(s)+\alpha_{s}^{*\top}\lambda_{s}-\tfrac{1}{2}\left|\alpha_{s}^{*}\right|^{2}\Big)\,ds+\int_{0}^{t}\alpha_{s}^{*\top}\,dB_{s}\right).$ (5.89)
@@
$\hskip 102.43008pt=\exp\Big(\gamma\int_{0}^{T}r(s)ds\Big)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}ds\Big)\Big],$
@@
$F_{s}^{\alpha^{*}}=\gamma\big(\alpha_{s}^{*\top}\lambda_{s}+\frac{\gamma-1}{2}\left|\alpha_{s}^{*}\right|^{2}\big)=\frac{\gamma}{2(1-\gamma)}\big(\left|\lambda_{s}\right|^{2}-\left|\Sigma\Lambda_{s}\right|^{2}\big)=\frac{\gamma}{2(1-\gamma)}\sum_{i=1}^{d}\Big(\theta_{i}^{2}-\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V^{i}_{s}$
@@
5.2 Proofs of Proposition 3.9 and Theorems 3.10 and 3.12
[…]
$\Gamma_{t}^{1-\rho^{2}}=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\Big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\left|\lambda_{s}\right|^{2}ds\Big)|\mathcal{F}_{t}\Big]=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}ds\Big)\mid\mathcal{F}_{t}\Big],$ (5.97)
$\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}-\frac{1-\rho^{2}}{2}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}ds\Big)\mid\mathcal{F}_{t}\Big]=\exp\Big(\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{1-\rho^{2}}{2}\theta_{i}^{2}+\tilde{F}_{i}(s,\tilde{\psi}(T-s))\big)\tilde{g}^{i}_{t}(s)ds\Big)$ (5.98)
[…]
$\tilde{\psi}^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\big(-\frac{\theta_{i}^{2}(1-\rho^{2})}{2}+\tilde{F}_{i}(T-s,\tilde{\psi}(s))\big)ds,\quad i=1,\ldots,d.$ (5.99)
where we used for the last identity the fact that $\left|\Sigma\Lambda_{t}\right|^{2}=\;\rho^{2}\sum_{i=1}^{d}\left(\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\right)^{2}V^{i}_{t},\;\text{and}\;\left|\lambda_{t}\right|^{2}=\;\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{t}.$ As $0<\Gamma_{t}\leq 1$, $\mathbb{P}-a.s.$, we have that $\Gamma\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})$. As for $\Lambda$, it is clear that it belongs to $L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ since $\varsigma$ and $\psi$ are bounded, $\psi$ is continuous thus bounded and $\mathbb{E}\Big[\int_{0}^{T}\sum_{i=1}^{d}V^{i}_{s}ds\Big]<\infty$ by (2.17). Consequently, $\left(\Gamma,\Lambda\right)\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. This complete the Proof $\Box$
Consequently, by Itô’s product rule, one may write
where the drift factor takes the form:
@@
$\mathbb{E}_{x_{0},V_{0}}\Big[U(X_{T}^{\alpha})\Big]$ $\,=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}\big(-\gamma e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\lambda_{s}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}\right|^{2}\big)ds\Big)\Big],$
$=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}^{\alpha}ds\Big)\Big],$
@@
Noting by a change of variables that
$\sum_{j=1}^{d}\psi^{j}(T-s)(DV_{s})_{j}=\sum_{j=1}^{d}\psi^{j}(T-s)\sum_{i=1}^{d}D_{ji}V_{s}^{i}=\sum_{i=1}^{d}V_{s}^{i}\sum_{j=1}^{d}D_{ji}\psi^{j}(T-s)=\sum_{i=1}^{d}(D^{\top}\psi)_{i}(T-s)V_{s}^{i},$
one gets using Equation (3.69)

### 06 Apr 2026 v2

@@ 5 Proofs of the main results
${Z}_{t}:=\mathcal{E}(\int_{0}^{t}g(s)\alpha_{s}^{\top}\,dB_{s})=\exp\Big(\int_{0}^{t}g(s)\alpha_{s}^{\top}\,dB_{s}-\frac{1}{2}\int_{0}^{t}g^{2}(s)\left|\alpha_{s}\right|^{2}ds\Big),\quad t\in[0,T].\vskip-5.69046pt$ (5.86)
@@ 2. $(b)$
We provide a proof of this classical result in Appendix A.3 for the reader’s convenience.
@@
$X_{t}^{\alpha^{*}}=x_{0}\exp\!\Big(\int_{0}^{t}\Big(r(s)+\alpha_{s}^{*\top}\lambda_{s}-\tfrac{1}{2}|\alpha_{s}^{*}|^{2}\Big)\,ds+\int_{0}^{t}\alpha_{s}^{*\top}\,dB_{s}\Big).\vskip-2.84544pt$ (5.91)
@@
$\hskip 102.43008pt=\exp\Big(\gamma\int_{0}^{T}r(s)ds\Big)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}ds\Big)\Big]$
@@
$F_{s}^{\alpha^{*}}=\gamma\big(\alpha_{s}^{*\top}\lambda_{s}+\frac{\gamma-1}{2}\left|\alpha_{s}^{*}\right|^{2}\big)=\frac{\gamma}{2(1-\gamma)}\big(\left|\lambda_{s}\right|^{2}-\left|\Sigma\Lambda_{s}\right|^{2}\big)=\frac{\gamma}{2(1-\gamma)}\sum_{i=1}^{d}\Big(\theta_{i}^{2}-\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V^{i}_{s}\vskip-5.69046pt$
@@
$\mathbb{E}_{x_{0},V_{0}}\Big[U(X_{T}^{\alpha})\Big]$ $\,=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}\big(-\gamma e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\lambda_{s}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}\right|^{2}\big)ds\Big)\Big]$
$=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}^{\alpha}ds\Big)\Big]$
@@
Recalling the change of variable in (5.98), one gets using Equation (3.70)
@@
Proof of Theorem 3.13: From the dynamic of the controlled wealth process given in Equation (3.31) and using the utility function $U(x):=\log(x)$, we have
$U(X_{T}^{\alpha}):=\log(X_{T}^{\alpha})=\log(x_{0})+\int_{0}^{T}\big(r(s)+\alpha_{s}^{\top}\lambda_{s}-\tfrac{1}{2}\left|\alpha_{s}\right|^{2}\big)\,ds+\int_{0}^{T}\alpha_{s}^{\top}\,dB_{s}.$ (5.108)
We consider the process $J_{t}^{\alpha}:=\log(x_{0})+\int_{0}^{t}\big(r(s)+\alpha_{s}^{\top}\lambda_{s}-\tfrac{1}{2}\left|\alpha_{s}\right|^{2}\big)\,ds+\int_{0}^{t}\alpha_{s}^{\top}\,dB_{s}+\Gamma_{t}$ where the pair $(\Gamma,\Lambda)$ satisfies a rather simple backward stochastic differential equation (BSDE) under $\mathbb{P}$ with a driver $f:[0,T]\times\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$ of the form:
$\left\{\begin{array}[]{ccl}d\Gamma_{t}&=&\;-f(t,\Gamma_{t},\Lambda_{t})dt+\Lambda_{t}^{\top}dW_{t},\\ \Gamma_{T}&=&0.\end{array}\right$. (5.109)
In these terms we are bound to choose a function $f$ for which $J_{t}^{\alpha}$ is a supermartingale for all $\alpha\in\mathcal{A}$ and there exists a $\alpha^{*}\in\mathcal{A}$ such that $J_{t}^{\alpha^{*}}$ is a martingale. We assume that $(\Gamma,\Lambda)\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R}^{*}_{+})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. Note that $J_{t}^{\alpha}$ satisfies conditions $1$ and $2$ of Definition 3.2 since $\Gamma_{T}=0$ and $\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$. Applying Itô’s formula on $J_{t}^{\alpha}$ yields $dJ_{t}^{\alpha}=\big(r(t)+\alpha_{t}^{\top}\lambda_{t}-\tfrac{1}{2}\left|\alpha_{t}\right|^{2}-f(t,\Gamma_{t},\Lambda_{t})\big)\,dt+\alpha_{t}^{\top}\,dB_{t}+\Lambda_{t}^{\top}dW_{t}$. In order for the third condition to be fulfilled by $J_{t}^{\alpha}$, we may have that $r(t)+\alpha_{t}^{\top}\lambda_{t}-\tfrac{1}{2}\left|\alpha_{t}\right|^{2}\leq f(t,\Gamma_{t},\Lambda_{t}),$ for all $t\in[0,T]$ and $\alpha\in\mathcal{A}$. The pointwise maximization yields $\alpha^{*}_{t}:=\lambda_{t}$ for every $t\in[0,T]$ and $dM_{t}:=\lambda_{t}^{\top}\,dB_{t}+\Lambda_{t}^{\top}dW_{t}$ is a true martingale by condition (3.27)( since Assumption 3.1 holds). The admissibility of $\alpha^{*}$ is straightforward thanks to the same assumptions. The corresponding value function is determined simply by the initial value of the supermartingale $J_{t}^{\alpha}$ at $\alpha^{*}$. Subtituting $\alpha^{*}$ into (5.108) and taking the expectation yields:
$\sup_{\alpha(\cdot)\in\mathcal{A}}\mathbb{E}\Big[\log(X_{T}^{\alpha})\Big]=\log(x_{0})+\mathbb{E}\Big[\int_{0}^{T}\big(r(s)+\tfrac{1}{2}\left|\lambda_{s}\right|^{2}\big)\,ds\Big]=\log(x_{0})+\int_{0}^{T}r(s)\,ds+\tfrac{1}{2}\int_{0}^{T}\sum_{i=1}^{d}\theta_{i}^{2}\mathbb{E}\Big[V^{i}_{s}\Big]\,ds$
This completes the proof, calling upon Proposition 2.4, and we are done. $\Box$
@@
Appendix A Supplementary materials and Proofs.
A.1 Proofs of Proposition 3.4 and Theorem 3.5
Proof of Proposition 3.4 : We write for  $0\leq t\leq T$:
$\Gamma_{t}^{\frac{1}{\delta}}=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}\frac{\gamma}{\delta}\big(r(s)+\frac{\left|\lambda_{s}\right|^{2}}{2(1-\gamma)}\big)ds\Big)\mid\mathcal{F}_{t}\Big]=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}\frac{\gamma}{\delta}\big(r(s)+\frac{1}{2(1-\gamma)}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}\big)ds\Big)\mid\mathcal{F}_{t}\Big]$ (A.110)
which ensures that $\Gamma_{t}>0$ $\mathbb{P}-a.s.$, since $V_{t}$ is non-negative ($V\in\mathbb{R}^{d}_{+}$), $r(t)>0$ is deterministic, and $1-\gamma\leq\delta\leq 1$. We then have in view of (A.110) that there exists some positive constant $m>0$ such that $\Gamma_{t}\geq m>0$ for every $t\in[0,T]$. An application of the exponential-affine transform formula (Gnabeyeu, 2026, Theorem A.4.) with $\mathcal{M}\ni m(\,\mathrm{d}s):=\frac{\gamma}{2\delta(1-\gamma)}\theta\odot\theta\,{\rm Leb}_{d}(\,\mathrm{d}s)$ (where $\odot$ denote the Hadamard (pointwise or component-wise) product) yields:
[…]
where we used for the last identity the fact that
where the drift factor takes the form:
$\hskip 142.26378pt\forall\,t\in[0,T]\quad J_{t}^{\alpha}=\frac{\Gamma_{0}x_{0}^{\gamma}}{\gamma}e^{\int_{0}^{t}D_{s}(\alpha_{s})\,ds}F_{t}^{\alpha}\quad\text{where}$ (A.111)
A.2 Proofs of Proposition 3.9 and Theorem 3.10
[…]
$\Gamma_{t}^{1-\rho^{2}}=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\Big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\left|\lambda_{s}\right|^{2}ds\Big)|\mathcal{F}_{t}\Big]=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}ds\Big)\mid\mathcal{F}_{t}\Big],$ (A.112)
$\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}-\frac{1-\rho^{2}}{2}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}ds\Big)\mid\mathcal{F}_{t}\Big]=\exp\Big(\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{1-\rho^{2}}{2}\theta_{i}^{2}+\tilde{F}_{i}(s,\tilde{\psi}(T-s))\big)\tilde{g}^{i}_{t}(s)ds\Big)$ (A.113)
[…]
$\tilde{\psi}^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\big(-\frac{\theta_{i}^{2}(1-\rho^{2})}{2}+\tilde{F}_{i}(T-s,\tilde{\psi}(s))\big)ds,\quad i=1,\ldots,d.$ (A.114)
where we used for the last identity the fact that
$\left|\Sigma\Lambda_{t}\right|^{2}=\;\rho^{2}\sum_{i=1}^{d}\left(\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\right)^{2}V^{i}_{t},\quad\text{and}\quad\left|\lambda_{t}\right|^{2}=\;\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{t}$
As $0<\Gamma_{t}\leq 1$, $\mathbb{P}-a.s.$, we have that $\Gamma\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})$. As for $\Lambda$, it is clear that it belongs to $L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ since $\varsigma$ and $\psi$ are bounded, $\psi$ is continuous thus bounded and $\mathbb{E}\Big[\int_{0}^{T}\sum_{i=1}^{d}V^{i}_{s}ds\Big]<\infty$ by (2.17). Consequently, $\left(\Gamma,\Lambda\right)\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. This complete the Proof $\Box$
Consequently, by Itô’s product rule, one may write
where the drift factor takes the form:
A.3 Proof of Lemma 5.3
The first claim in (5.89) follows by the associativity of the convolution and applying the fundamental theorem of calculus. For the second claim, one could employ the calculus of convolutions and resolvents. However, we instead use the Laplace transform to deduce that: